Henrik's Options. Assume Henrik writes a call option on euros with a strike price of ​$1.2500​/euro at a premium of 3.80cents per euro ​($0.0380​/euro​) and with an expiration date three months from now. The option is for euro100 comma 000. Calculate​ Henrik's profit or loss should he exercise before maturity at a time when the euro is traded spot at strike prices beginning at ​$1.10​/euro​, rising to ​$1.34​/euro in increments of ​$0.04. The profit or loss should Henrik exercise before maturity at a time when the euro is traded spot at ​$1.10​/euro is ​$ nothing. ​(Round to the nearest cent and indicate a loss by using a negative​ sign.)
Answer #1

1.
=100000*(MAX(1.1-1.25,0)-0.038)=-3800

2.
=100000*(MAX(1.14-1.25,0)-0.038)=-3800

3.
=100000*(MAX(1.18-1.25,0)-0.038)=-3800

4.
=100000*(MAX(1.22-1.25,0)-0.038)=-3800

5.
=100000*(MAX(1.26-1.25,0)-0.038)=-2800

6
=100000*(MAX(1.3-1.25,0)-0.038)=1200

7
=100000*(MAX(1.34-1.25,0)-0.038)=5200.00000000001